After a 1 year investment you receive 8.5% interest (nominal) from your bank. However, looking at how prices have changed, you soon realize that the real rate of interest was actually 1.6%. How much was inflation during that year?
Answer #1

Based on Fischer relation,

(1 + Nominal rate) = (1 + Real rate) * (1 + Inflation)

(1 + 8.5%) = (1 + 1.6%) * (1 + Inflation)

1.085 = 1.016 * (1 + Inflation)

1 + Inflation = 1.085/1.016 = 1.0679

Inflation = 0.0679

Inflation = 6.79%